OpenStudy (anonymous):

GIVING MEDALS The coefficient of kinetic friction between a couch and the floor is 0.4. If the couch has a mass of 35 kg and you push it with a force of 200 N, what is the net force on the couch as it slides?

OpenStudy (zephyr141):

\[\sum_{}^{}F _{x}:ma=-f _{k}+200=0\]\[\sum_{}^{}F_{y}:n-mg=0 \rightarrow n=mg\] \[f_{0}=\mu*n=\mu*m*g\] so \[\sum_{}^{}F=200-(0.4)(35)(9.8)\] i think... i hope i did this right.